OpenStudy (anonymous):

An archer shoots an arrow with an initial velocity of 21 m/s straight up from his bow. He quickly reloads and shoots another arrow in the same way 3.0 s later. At what time and height do the arrows meet?

OpenStudy (anonymous):

explain how to solve this pls

OpenStudy (anonymous):

mmh okay

OpenStudy (anonymous):

msg me once u r here

OpenStudy (anonymous):

@aajugdar how do you solve this problem?

OpenStudy (anonymous):



OpenStudy (anonymous):

Vi = 21 m/s

OpenStudy (anonymous):

first we have to find the max time to reach at maximum height for the first arrow

OpenStudy (anonymous):

equation of motion is... u = v + at where u is final velocity

OpenStudy (anonymous):

u = 0 v = 21 m/s a = -g ( while arrow moving upward).

OpenStudy (anonymous):

so we have 0 = (21 m/s) - gt ===> g = 9.8 m/ sec^2

OpenStudy (anonymous):

then what to do next?

OpenStudy (anonymous):

so you have g= 9.8 put the value and you get time

OpenStudy (anonymous):

which is 2.14

OpenStudy (anonymous):

2.14 is the time when it reaches the top so when its to the ground,then total time would be 4.28 s

OpenStudy (anonymous):

now see another arrow is thrown at same velocity

OpenStudy (anonymous):

after 3 seconds

OpenStudy (anonymous):

look at this

OpenStudy (anonymous):

so it will meet first arrow when first arrow is coming down

OpenStudy (anonymous):

okay, so what will be the time when they meet? and how do i solve for the height when the 2 arrows meet?

OpenStudy (anonymous):

you will hv to find displacement

OpenStudy (anonymous):

so you can see the formula for displacement in that link

OpenStudy (anonymous):

s = ut +0.5at^2

OpenStudy (anonymous):

u is initial velocity which is 21 t is time which is 2.14

OpenStudy (anonymous):

displacement of the arrow will be the height it reaches

OpenStudy (anonymous):

now you know that 1st arrow is falling when 2nd is fired right?

OpenStudy (anonymous):

so it starts falling after 2.14 secs and 2nd arrow is fired after 3 seconds so it will be falling for 3-2.14 = 0.86 secs when 2nd arrow is fired

OpenStudy (anonymous):

So after falling for 0.85 seconds: s = 0 + 0.5(9.8)(0.85)^2 = 3.54m you understood this part?

OpenStudy (anonymous):

its displacement from top is 3.54 so first arrow's height will be 22.5 - 3.54 = 18.960m

OpenStudy (anonymous):

what about the 2nd arrow's height?

OpenStudy (anonymous):

thats the height of first arrow when 2nd arrow is fired,so it will be 0 then

OpenStudy (anonymous):

ok. so at what height do they meet?

OpenStudy (anonymous):

before that you will have to find time and before that velocity of first arrow when its coming down

OpenStudy (anonymous):

Its velocity is: v = at = 9.8(0.85) =8.33m/s

OpenStudy (anonymous):

now final part is easy as the total distance between the two arrows just when arrow 2 is being fired is 18.96m, then the distance traveled by arrow 1 is 18.96-s, where s is the distance traveled by arrow 2.

OpenStudy (anonymous):

now u know the equation s= ut+0.5at^2

OpenStudy (anonymous):

u hv velocities

OpenStudy (anonymous):

all u hv to do is keep time same in both equations and put the values (18.96-s) = 8.33t + 4.6t^2.....arrow 1 s = 21t - 4.6t^2.......arrow 2 subbing in arrow 1 into arrow 2 EQN yields: 18.96 - (21t - 4.6t^2) = 8.33t + 4.6t^2 29.33t = 18.96

OpenStudy (anonymous):

t = 0.6464seconds This is the answer for time relative to the firing of arrow 2. relative to start of arrow 1's journey, time is 2.14 + 0.85 + 0.6464 = 3.636s

OpenStudy (anonymous):

now you have time and velocity of first arrow if u find dispacement of first arrow it will be the height

OpenStudy (anonymous):

the height where they meet

OpenStudy (anonymous):

so s = ut +0.5at^2 = 21(0.6464) - 4.6(0.6464)^2 = 11.65m from the ground

OpenStudy (anonymous):

thats the height

OpenStudy (anonymous):

ok. @aajugdar thank you so much!

OpenStudy (anonymous):

yaw

OpenStudy (anonymous):

My net connection was get off.... :(

OpenStudy (anonymous):

@aajugdar shouldn't be the time 0.86 instead of 0.85?

OpenStudy (anonymous):

yes but wont make much of a difference